Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 17:05 Wednesday 20 January 2021 (15798) Print this report
BF ballast mass adjustment.

With Washimi-san

See pictures in the album BS Remedying work after O3.

Today we only did preparation work as follows:

  • I moved the IM-Y close to zero using the IM yaw picomotor and kept the oplev in a suitable placearound zero yaw compensating with the new F0-Y stepper motor.
  • I measured TFs for the six DoF of the IM as health check; all were alright.
  • The temperature in the clean booth was 23.4 degrees.
  • Without fastening the ballast masses we added 2 X 31 grams on top of the BF, action which brought the SF keystone down to 226 um from approximately 500 um. Our aim is 212 um, so it's very close.
  • At the end we removed the masses because we didn't put them in a place where we could fasten them.

Tomorrow we will add the masses in the appropriate places and we will fasten them.

Comments to this report:
fabian.arellano - 17:36 Thursday 21 January 2021 (15806) Print this report

With Washimi-san

See pictures in the album BS Remedying work after O3.

We proceeded to add ballas mass.

  • See these related reports: klog 15651 and klog 15658.
  • See the conditions of the initial positions of the different stages and temperature in this screenshot and in my notes
  • Initial temperature: 23.1 degrees
  • Final temperature: 23.5 degrees.

We did two iterations:

  • First, we added 2 X 35 gr, in the -Y and +Y sides of BF top, and the F1 keystone went down to around 203 µm from 521 µm.
  • Later we removed two grams each side, this is, we added 2 X 33 gr in total, however, the keystone remained at around 203 µm due to the large temperature increase.


  • The temperature increase didn't allow us to realistically assess the effect of adding mass, but the F1 keystone is already reltively close to 212 µm, namely, it's about 16 µm below the goal at 23.4 degrees.
  • The screw fastening the masses in the +Y side was tightened by hand and for the one in the -Y side I gently used an Allen key.
  • There was no significant change in the readout of the OSEMs related to the addition of mass (V1, V2 and V3 changed due to temperature change), and the oplev changed an yaw a little bit, which was expected. See this picture for the condition at the time of writing.
  • I measured transfer functions of the IM and the system looks healthy.
  • See details of the procedure in the notebook pages here and here.
  • See the pictures in the album also.
tatsuki.washimi - 7:50 Friday 22 January 2021 (15809) Print this report

The temeprature of BS and SR2 were increased during our work and went back after the work.

We suspect it caused by the LED lights and/ or KOACH filters.

Images attached to this comment
shinji.miyoki - 8:24 Friday 22 January 2021 (15810) Print this report

One KOACH corresponds to ~ 200 W heat generator.

Search Help